OpenStudy (anonymous):

How do you factor x^2 + 7x + 6 and 49x^2 - 16? Pleasee help :)

OpenStudy (anonymous):

Remember to start out with listing what the constant can be divided by: for a) 6, 1, 3, 2 and for b) 4, 4, 1, 16

OpenStudy (anonymous):

Now for a), look at the middle factor. Are there any two numbers that multiply to six and add to seven?

OpenStudy (anonymous):

Once you find that, just put it in (x+a)(x+b) form and you're done.

OpenStudy (anonymous):

The second one's a bit more complicated. There is no middle term. This is of course because it is 0. So, you need to find two numbers that multiply to -16 and add to 0.

OpenStudy (anonymous):

As well, with b), you must take in account the coefficient of x^2, which is 49.

OpenStudy (anonymous):

With this type of question, the rule of thumb is that you use the square root of the coeffienct or constant for your factoring; in this case, 7 and 4.

OpenStudy (anonymous):

So the answer would be (7x+4)(7x-4)

OpenStudy (anonymous):

(for b)

OpenStudy (whpalmer4):

You should train yourself to look at any expression of the form <something> - <something else> and ask yourself if they could both be squares. The difference of squares is readily factored: \[a^2- b^2 = (a+b)(a-b)\] Here we have \(49x^2-16\) so \(a^2=49x^a\) and \(b^2=16\) That gives us \(a = 7x\) and \(b = 4\), so the factoring is \[49x^2-16 = (a+b)(a-b) = (7x+4)(7x-4)\]